OpenStudy (anonymous):

Coach Johnson is 10 years older than Tim. When Tim started football 3 years ago, Coach Johnson was twice his age. How old is Tim now?

OpenStudy (johnweldon1993):

Well lets let C = coach and T = Tim so C = T + 10 and we can set up the equation from the question as C-3 = 2(T-3) and since we know that C = T + 10...we can replace C with that... (T + 10) - 3 = 2(T-3) and simplify this.. T + 10 - 3 = 2T - 6 T + 7 = 2T - 6 subtract T from both sides: T + 7 = 2T - 6 -T -T 7 = T - 6 And finally add 6 to both sides to solve for Tims age... T = ???

OpenStudy (anonymous):

t = 13?

OpenStudy (johnweldon1993):

Right...now lets check it If Tim is now 13 And the coach is 10 years older than him...that would make the coach 23... and if 3 years ago the coach was double Tim's age...that means 2(13 - 3) = (23 - 3) 20 = 20 so yes 13 is Tim's age now, and coach is 23; so three years ago they were 10 & 20 Make sense? So yes Tim is 13 years old

OpenStudy (anonymous):

yes! thank you!

OpenStudy (johnweldon1993):

No problem! :)